OpenStudy (anonymous):

What is the temperature change for a substance if 198 g releases 0.234 kJ of heat and has a specific heat of 0.360 J/g°C? And why?

OpenStudy (aaronq):

use the equation: \(q=m*C_p*\Delta T\)

OpenStudy (anonymous):

it says the answer is 3.28 how do I get that answer

OpenStudy (aaronq):

plug your variables into the equation

OpenStudy (anonymous):

can you lay it out for me with the variables plugged in

OpenStudy (aaronq):

nope, cuz then i'd be doing it for you and you wouldn't learn anything. you should try it, i'll help you correct it

OpenStudy (anonymous):

all I keep getting is 16.67

OpenStudy (aaronq):

show me what you've done and i'll help you correct it

OpenStudy (anonymous):

I did 198*0.234*.360

OpenStudy (aaronq):

you're looking for the change in temperature, \(\Delta T\). do you know what the other variables in this "\(q=m*C_p*\Delta T\)" are?

OpenStudy (anonymous):

I know m is 198 and Cp is 0.360 so does that mean q is .234?

OpenStudy (aaronq):

exactly. do you know what the sign of q (positive or negative) should be when the system is releasing energy?

OpenStudy (anonymous):

I do not, I am sorry I know this must be frustrating but I was not supposed to be in Chemistry in the first place, let alone honors

OpenStudy (aaronq):

no worries, it's pretty easy if you think about it. It always helps to know the definitions of everything. So when a system is releasing energy, it's called an exothermic process, q is negative. Conversely, when a system is absorbing energy, it's called an endothermic process, q is positive.

OpenStudy (aaronq):

can you plug the variables into the equation with the correct signs now?

OpenStudy (anonymous):

-0.234=198*0.360* ΔT

OpenStudy (aaronq):

yep that's good. except 0.234 kJ and 0.360 J/g*C need to be in the same units. 1 kJ = 1000 J so 0.234 kJ * 1000J/kJ= 234 J -234=198*0.360* ΔT

OpenStudy (aaronq):

this will give you the right answer.

OpenStudy (anonymous):

I am still not getting it, how do I solve with the -.234 on the left side

OpenStudy (aaronq):

remember it's not -0.234 kJ it's -234.0 J, units need to be the same. solve it like any algebraic equation: -234=198*0.360* ΔT isolate ΔT=\(\dfrac{-234\;J}{198\;g*0.360\;J/g*^oC}\)

OpenStudy (anonymous):

thank you so much that really really helped

OpenStudy (aaronq):

no problem ! i hope you understood it well enough to do it by yourself next time